I like the gather environment, useful when you don’t want things aligned on the equal sign.

f(x) = x + |x + 3| − |x − 3|  (1)
−3 ≤ x < 3  (2)

I like the split environment, which allows nice formatting of equations too long to fit on one line.

\begin{aligned}f(x) & = f(x) + f′(a)(x − a) + \frac{f′′(a)}{2!}(x − a)^{2} + \frac{f′′′(a)}{3!} + ⋯ \\ & + \frac{f^{(n)}(a)}{n!}(x − a)^{n} + \frac{f^{(n+1)}(ξ)}{(n + 1)!}(x − ξ)^{n+1} \end{aligned}  (3)